Please see the problem below, please be sure to read the numbers and question carefully and be sure to check for all unit conversions and double check your answer. Thank you!

D | Question 3 18 pts A charge is located at the origin, another charge is located somewhere on the +y axis, another charge is located somewhere on the -y axis, and a fourth is located somewhere on the +x axis. The angle of the total electric force on the charge at the origin is 38 degrees counter-clockwise from the +x axis. Then, the charge at the origin changes sign and its magnitude (absolute value of its charge) increases by a factor of 2.6. All other charges do not move or change their signs or magnitudes. What is the angle of the total electric force on the charge at the origin now measured in degrees counter-clockwise from the tx axis?
Answer #1

when the charge at origin is changed 3 21 21 CO 3 2-

when the charge is made 2.6 times , each force also becomes 2.6 times and hence that 2.6 factor cancel out in the ratio of the forces while calculating the angle.

from case 1 ,

tan38 = (F2 - F3)/F4

from case 2 ,

tanheta = (F2 - F3)/F4

hence

heta = 38

angle from the positive x-axis is given as

phi = 270 - heta

phi = 270 - 38

phi = 232 deg